Order: Heparin 7,500 units, SubQ, q8hrs Drug Available: Heparin 5,000 units/1mL How many units will the...

  1. Order: Heparin 7,500 units, SubQ, q8hrs Drug Available: Heparin 5,000 units/1mL
  1. How many units will the patient recieve per day?
Answer #1

Given

Order:HeparinDose:7500units

Route:subQ

Frequency:Thrice a day

Available:5000units/ml

When 5000 units=1ml

Then 7500units=?ml

Cross multiply and divide

=7500×1÷5000

=7500÷5000

=1.5 has to be given per dose

If one dose:7500units then 3 doses are 3×7500=22500

Ans:22,500units per day